OpenStudy (anonymous):

The half-life of cesium-137 is 30 years. Suppose we have a 120 mg sample. (a) Find the mass that remains after t years.

OpenStudy (anonymous):

These systems die off in proportion to themselves. So the rate of decay is given by \[\frac{dx}{dt}=-kx\] where k is positive. The negative sign ensures the rate is a decaying one. This is separable, so\[\frac{dx}{x}=-k dt \rightarrow \int\limits_{}{}\frac{dx}{x}=\int\limits_{}{}-kdt=\log x =-\]

OpenStudy (anonymous):

log x = -kt +c

OpenStudy (anonymous):

k= (ln 2) / 30

OpenStudy (anonymous):

Exponentiate both sides to get,\[x(t)=x_0e^{-kt}\]

OpenStudy (anonymous):

Now, at time t=0, you have 120mg, so\[x(0)=120mg=x_0e^0=x_0\]

OpenStudy (anonymous):

x(t) = 120 x e^((ln2)/30)

OpenStudy (anonymous):

x(t) = 120 x e^-((ln2)/30)

OpenStudy (anonymous):

Also, after 30 years, 1/2 the mass remains, so\[x(30)=60=120 e^{-30k} \rightarrow \frac{1}{2}=e^{-30k} \rightarrow k=-\frac{1}{30}\log \frac{1}{2}\]

OpenStudy (anonymous):

i.e. \[k=\frac{\log 2}{30}\]So,\[x(t)=120e^{-\frac{\log 2}{30}t}\]

OpenStudy (anonymous):

Thanks suzi20 and lokisan

OpenStudy (anonymous):

Welcome. Fan us!